# Computation and Communication Co-scheduling for Timely Multi-Task Inference at the Wireless Edge

### v1

@@ Appendix B Proof of Theorem 1
Because $p_{m,j}(\delta)$ is bounded, there exist finite constants $\bar{p}_{h}$ and $\bar{p}_{l}$ such that $\bar{p}_{l}\leq w_{m,j}p_{m,j}(\delta)\leq\bar{p}_{h}$. Let $\bar{p}_{opt}(T)$ and $\bar{p}_{{\text{MGF}}}(T)$ denote the discounted sum of inference errors under an optimal policy to (III-A)-(6) and the MGF policy, respectively, truncated at time $T$. Then, we have
@@
${\leq}~{}\bar{p}_{{\text{MGF}}}(T)-\bar{p}(\boldsymbol{\lambda^{*}},\boldsymbol{\mu^{*}};T)+\frac{\gamma^{T+1}(\bar{p}_{h}-\bar{p}_{l})}{1-\gamma},$ (27)
where the inequality holds because the penalty functions are bounded and the weak duality $\bar{p}(\boldsymbol{\lambda^{*}},\boldsymbol{\mu^{*}};T)\leq\bar{p}_{opt}(T)$.
@@
Similar to [12, corollary 4.4], we can show the following Lemma:
@@
$\bar{p}_{{\text{MGF}}}(T)-\bar{p}(\boldsymbol{\lambda^{*}},\boldsymbol{\mu^{*}};T)\leq\frac{\gamma(\bar{p}_{h}-\bar{p}_{l})max_{t}B_{t}}{(1-\gamma)^{3}\sum_{m=1}^{M}k_{m}}.$ (29)
@@
$\bar{p}_{{\text{MGF}}}(T)-\bar{p}(\boldsymbol{\lambda^{*}},\boldsymbol{\mu^{*}};T)\leq\frac{2(\bar{p}_{h}-\bar{p}_{l})\gamma\sum_{m=1}^{M}\sqrt{k_{m}}}{(1-\gamma)^{3}\sum_{m=1}^{M}k_{m}}$
$\!\!\!\!\!\!\leq\frac{2M(\bar{p}_{h}-\bar{p}_{l})\gamma\sum_{m=1}^{M}\sqrt{k_{m}}}{(1-\gamma)^{3}(\sum_{m=1}^{M}\sqrt{k_{m}})^{2}}\leq\frac{2M(\bar{p}_{h}-\bar{p}_{l})\gamma}{(1-\gamma)^{3}\sum_{m=1}^{M}\sqrt{k_{m}}},$ (30)

### v2

@@ Appendix B Proof of Theorem 1
Because $p_{m,j}(\delta)$ is bounded, there exist finite constants $\bar{p}_{h}$ and $\bar{p}_{l}$ such that $\bar{p}_{l}\leq w_{m,j}p_{m,j}(\delta)\leq\bar{p}_{h}$. Let $\bar{p}_{opt}(T)$ and $\bar{p}_{{\text{MGF}}}(T)$ denote the discounted sum of inference errors under an optimal policy to (4)-(6) and the MGF policy, respectively, truncated at time $T$. Then, we have
@@
${\leq}~\bar{p}_{{\text{MGF}}}(T)-\bar{p}_{1}(\boldsymbol{\lambda}^{*}_{1:T},\boldsymbol{\mu}^{*}_{1:T})+\frac{\gamma^{T+1}(\bar{p}_{h}-\bar{p}_{l})}{1-\gamma},$ (36)
where the inequality holds because the penalty functions are bounded and the weak duality $\bar{p}_{1}(\boldsymbol{\lambda}^{*}_{1:T},\boldsymbol{\mu}^{*}_{1:T})\leq\bar{p}_{opt}(T)$.
@@
Similar to [13, corollary 4.4], we can show the following Lemma:
@@
$\bar{p}_{{\text{MGF}}}(T)-\bar{p}_{1}(\boldsymbol{\lambda}^{*}_{1:T},\boldsymbol{\mu}^{*}_{1:T})\leq\frac{\gamma(\bar{p}_{h}-\bar{p}_{l})max_{t}B_{t}}{(1-\gamma)^{3}\sum_{m=1}^{M}k_{m}}.$ (38)
@@
$\bar{p}_{{\text{MGF}}}(T)-\bar{p}_{1}(\boldsymbol{\lambda}^{*}_{1:T},\boldsymbol{\mu}^{*}_{1:T})$ $\leq\frac{2(\bar{p}_{h}-\bar{p}_{l})\gamma\sum_{m=1}^{M}\sqrt{k_{m}}}{(1-\gamma)^{3}\sum_{m=1}^{M}k_{m}}$
$\leq\frac{2M(\bar{p}_{h}-\bar{p}_{l})\gamma\sum_{m=1}^{M}\sqrt{k_{m}}}{(1-\gamma)^{3}(\sum_{m=1}^{M}\sqrt{k_{m}})^{2}}$
$\leq\frac{2M(\bar{p}_{h}-\bar{p}_{l})\gamma}{(1-\gamma)^{3}\sum_{m=1}^{M}\sqrt{k_{m}}},$ (39)